Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by HarvardDreamin » Fri Mar 21, 2008 9:00 am
Q2) A shipment of 8 television sets contains 2 black-and-white sets and 6color sets. If 2 television sets are to be chosen at random from this shipment, what is the probability that at least 1 of the 2 sets chosen will be a black-and-white set?

A. 1/7 B. ¼ C. 5/14 D. 11/28 E. 13/26
ON MY WAY TO HBS......
Join the discussion
Source: — Problem Solving |

by xilef » Fri Mar 21, 2008 10:10 am
probability that at least 1 of the 2 sets chosen will be a black-and-white set means that either one or both can be black-and-white.

First one yes, second no - 1/4 * 6/7 = 6/28
First one no, second yes - 3/4 * 2/7 = 6/28
Both are - 1/4 * 1/7 = 1/28

Now add all three scenarios:

6/28+6/28+1/28=13/28

Were the answers copied correctly?
Join the discussion

by musicdaemon » Fri Mar 21, 2008 10:45 am
i also think the ans is copied wrong,

i used the combinatorics to solve,

at least 1 black & white set is to be selected, i.e either one or both the sets are B&W.

Suppose, case is to select none of the B&W sets,

n(E)= number of events of two color sets being selected
= 6C2 = 15
n(S)= number of events of selecting any two sets
= 8C2 = 28

probability of selecting exclusively two color sets,

n(P) = (n(E))/(n(S))= 15/28

let n(U)=any two sets are selected=1

Therefore, for at least one B&W set to be selected

n(P') = 1-n(P)
= 1-(15/28) = 13/28............................... Ans.(xilef already solved)
Let the Game begin
Join the discussion

by HarvardDreamin » Fri Mar 21, 2008 11:41 am
Had the same query when I first solved it and thought I was incorrect.
The answers are exactly the same as in the question. I got it from the Sets package of questions. Im guessing either we're all missing something or the question is fundamentally flawed.
ON MY WAY TO HBS......
Join the discussion

by beeparoo » Sun May 11, 2008 5:36 pm
xilef wrote:probability that at least 1 of the 2 sets chosen will be a black-and-white set means that either one or both can be black-and-white.

First one yes, second no - 1/4 * 6/7 = 6/28
First one no, second yes - 3/4 * 2/7 = 6/28
Both are - 1/4 * 1/7 = 1/28

Now add all three scenarios:

6/28+6/28+1/28=13/28

Were the answers copied correctly?
What is the Answer according to the Set; did the Set come with an answer at all?

To me, the question stem does not distinguish between the order of selecting a B&W set first, or, second -- only that it would be selected AT LEAST once. While I agree with your approach in that both B&W sets may be selected too, I am not fully convinced that you need to account for the following:

"First one yes, second no"
"First one no, second yes"

Wouldn't the probability of one B&W and one colour (in no particular order) suffice?

Using this approach leads to answer B.

What do you think?
Join the discussion